OpenStudy (anonymous):

1. - 5(2x+2) - 10x - 17 (1 point) xis greater than or equal to7 xis greater than or equal to19 all real numbers no solution

OpenStudy (mayankdevnani):

WELCOME TO OPEN STUDY. @blobster2

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

-5(2x+2)-10x-17 -10x+(-10)-10x-17 = -20x-27

OpenStudy (mayankdevnani):

\[\huge -5(2x+2)=-5 \times 2x + (-5) \times 2\]

OpenStudy (mayankdevnani):

can you solve this? @blobster2

OpenStudy (anonymous):

no i tried

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

Mathway.com @blobster2

OpenStudy (mayankdevnani):

@M.A.X.I.N.E please don't give direct answers..... it not our CODE OF CONDUCT.... first read CODE OF CONDUCT... and sorry for this rude behaviour but what can i do? I have to follow the guidelines of OPEN STUDY.

OpenStudy (mayankdevnani):

i hope you understand my feelings. @M.A.X.I.N.E

OpenStudy (mayankdevnani):

can you solve it? @blobster2

OpenStudy (anonymous):

- 5(2x+2)is greater than or equal to - 10x - 17 thats the right equation

OpenStudy (mayankdevnani):

\[\huge -5(2x+2)=-10x-10\] that was correct ^^^

OpenStudy (mayankdevnani):

understood? @blobster2

OpenStudy (mayankdevnani):

then, \[\huge -10x-10+[-10x-17]\] open this bracket, \[\huge -10x-10-10x-17\] what we get after solving it? @blobster2